OpenStudy (anonymous):

PYTHAGOREAN THEOREM What is the length of the shortest side of a triangle that has vertices at (-3, -2), (1, 6), and (5, 3)? A.SQUARE ROOT OF 7UNITS B.9 UNITS C.5 UNITS D. 2 TIMES THE SQUARE ROOT OF 5 UNITS NEED HELP A.S.A.P

OpenStudy (anonymous):

 Sides are: \[\sqrt{(1--3)^2 + (6--2)^2} = \sqrt(16+64) = \sqrt80\ = 4\sqrt5\] \[\sqrt{(5-1)^2+(3-6)^2} = \sqrt(16+9) = 5\] \[\sqrt{(5--3)^2+(3--2)^2} = \sqrt{64+25} = \sqrt89\] Therefore clearly the shortest side is 5 units.

OpenStudy (anonymous):

THANKS